OpenStudy (vortish):

find the sume of the geometric series by using a formula 1-2+4-8+16-032+64

OpenStudy (zzr0ck3r):

sumation (-2)^k do you know the formula for geometric series?

OpenStudy (vortish):

no

OpenStudy (anonymous):

ok recall the formula

OpenStudy (anonymous):

\[\sum_{i=1}^{n}a \left( \frac{1-r^n}{1-r} \right)\]

OpenStudy (anonymous):

you have seven terms, and your first term =1

OpenStudy (anonymous):

by the way there is a simple way to do this by inducing the terms you provided let me show you

OpenStudy (anonymous):

I did the problem and here what i did and can you please tell me if I did it right please and thank you. 1-2+4-8+16-32+64= -1+4-8+16-32+64= 3-8+16-32+64= -5+16-32+64= 11-32+64= -21+64=43 Did i do it right??

OpenStudy (anonymous):

he needs to use a GS formula to solve that

OpenStudy (anonymous):

ok let me show you a simple way to solve this

OpenStudy (anonymous):

\[S_n = 1-2+4-8+16-32+64+...+(-2^n)\]

OpenStudy (anonymous):

so a =1 , r=-2, say n=3 the expected value for the sum would be Sn=3, let's check that

OpenStudy (anonymous):

recall \[\sum_{i=1}^{n}a \left( \frac{1-r^n}{1-r} \right)\]

OpenStudy (anonymous):

\[\sum_{i=1}^{3}(1) \left( \frac{1-(-2)^3}{1-(-2)} \right)=\sum_{i=1}^{3}\left( \frac{9}{3} \right)=3\]

OpenStudy (anonymous):

so the sum of the first 3 terms is three let's check it, we have 1-2+4=3, so our answer is correct

OpenStudy (anonymous):

now counting the number of terms you provided in the question, which is 7 terms that means now n=7

OpenStudy (anonymous):

\[\sum_{i=1}^{7}(1) \left( \frac{1-(-2)^7}{1-(-2)} \right)=\sum_{i=1}^{7}\left( \frac{129}{3} \right)=43\]

OpenStudy (anonymous):

let us check 1-2+4-8+16-32+64= 43

OpenStudy (anonymous):

This is your final solution, i hope it helped...